# On the Provable Suboptimality of Momentum SGD in Nonstationary Stochastic Optimization

### v2

@@ Abstract
While momentum-based acceleration has been studied extensively in deterministic optimization problems, its behavior in nonstationary environments—where the data distribution and optimal parameters drift over time—remains under-explored. We analyze the tracking performance of Stochastic Gradient Descent (SGD) and its momentum variants (Polyak Heavy-Ball and Nesterov) under uniform strong convexity and smoothness in varying step-size regimes. We derive finite-time, expectation and high-probability bounds for the tracking error, establishing a sharp decomposition into three components: a transient initialization term, a noise-induced variance term, and a drift-induced tracking lag. Crucially, our analysis uncovers a fundamental trade-off: while momentum can suppress gradient noise, it incurs an explicit penalty on the tracking capability. We show that momentum can substantially amplify drift-induced tracking error, with amplification that becomes unbounded as the momentum parameter approaches one, formalizing the intuition that satisfying "stale" gradients hinders adaptation to rapid regime shifts. Complementing these upper bounds, we establish minimax lower bounds for dynamic regret under gradient-variation constraints. These lower bounds prove that the inertia-induced penalty is not an artifact of analysis but an information-theoretic barrier: in drift-dominated regimes, momentum creates an unavoidable "inertia window" that fundamentally degrades performance. Collectively, these results provide a definitive theoretical grounding for the empirical instability of momentum in dynamic environments and delineate the precise regime boundaries where SGD provably outperforms its accelerated counterparts.
@@ 1 Introduction
Consider the optimization problem posed by a strongly convex objective function $G:\mathbb{R}^{d}\rightarrow\mathbb{R}$ defined as follows:
$\bm{\theta}^{\star}\in\operatorname*{arg\,min}_{\bm{\theta}\in\mathbb{R}^{d}}G(\bm{\theta}),\;\;\;G(\bm{\theta})=\mathbb{E}_{X\sim\Pi}[g(\bm{\theta},X)].$
Here $g(\bm{\theta},X)$ is a noise perturbed measurement of $G(\bm{\theta})$ and $X$ is a random variable sampled from a distribution $\Pi$. This is a classic formulation of stochastic optimization which serves as the theoretical and practical backbone of modern machine learning and has been explored across many domains [BB07, POW19, CLL+24, WS21, GCP+16]. Over the last two decades, finite-time guarantees for stochastic approximation methods have matured substantially [ACD+14, MB11, LXZ19, NJL+09], but the predominant assumption is stationarity of $\Pi$ i.e. the data are generated from a fixed distribution throughout the run of the algorithm. In many real systems—especially nonstationary time series—this assumption is violated due to regime changes, concept drift, or unmodeled exogenous effects and results in the data distribution evolving over time as $\Pi_{t}$. These settings naturally connect to classical stochastic tracking in signal processing [KY97, SAY03] and concept drift in online learning [HAZ16, CDH21], where the goal is not to converge to a single static solution, but to track an unknown, time-varying sequence of minimizers as closely as possible.
The primary algorithmic tool for this tracking task, provided that $g$ is differentiable, is Stochastic Gradient Descent (SGD) [RM51], especially when dealing with high-dimensional covariates $X_{t}$ and the number of training data points is large. SGD replaces the unavailable population gradient by noisy gradients computed from samples (or mini-batches). Due to its simplicity, scalability to big data, and robustness to noise and uncertainty, stochastic gradient methods have become popular in large-scale optimization, machine learning, and data mining applications [ZHA04, CBS14, SLJ+15].
Developing methods to accelerate the convergence of SGD has been studied extensively. Momentum methods, such as Polyak’s heavy-ball (HB) [POL63] and Nesterov acceleration (NAG) [NES83], are ubiquitous in deep learning and are widely believed to reduce gradient noise and accelerate training [SMD+13]. It is widely known for a variety of strongly convex and nonconvex settings that SGD with momentum provides faster convergence and better generalization compared to SGD without momentum [RAC+24, SI24, LGY20b, CM20]. Yet this deterministic intuition becomes fragile in genuinely nonstationary regimes: when the data distribution $\Pi_{t}$ drifts, past gradients may become systematically stale, so the same temporal averaging that suppresses noise can induce tracking lag and even instability, as observed empirically in online forecasting and suggested by prior steady-state analyses in the slow-adaptation regime.
@@
When does momentum help—and when does it provably hurt—in stochastic optimization under distribution shift? Concretely, can we (a) construct explicit drifting instances where Heavy-Ball/Nesterov incur an unavoidable inertia window and are uniformly worse than SGD, and (b) characterize the precise regime boundaries (in terms of drift, noise, conditioning, and momentum $\beta$) under which momentum improves variance reduction without sacrificing tracking accuracy?
@@
    Finite-time tracking guarantees and a concrete momentum penalty. We prove finite-time tracking bounds that cleanly separate three effects: forgetting of the initialization, an irreducible noise floor, and an irreducible tracking lag due to drift. Extending the analysis to Heavy-Ball and Nesterov, we obtain the same decomposition but show that momentum incurs an explicit penalty as $\beta$ tends to $1$: it becomes more sensitive to initialization, inflates the effective noise level, and requires a more conservative stepsize for stability. These effects are especially pronounced in ill-conditioned problems.
@@
    Drift-adaptive high-probability bounds that expose a new drift–noise coupling. Prior high-probability results often summarize nonstationarity by a single worst-case drift parameter, which cannot distinguish bursty regime shifts from persistent drift and ignores exponential forgetting. In contrast, our bounds are time-resolved: only recent drift meaningfully affects the guarantee. Moreover, the bounds reveal a drift–noise interaction mechanism: drift increases tracking mismatch, which in turn amplifies stochastic fluctuations. This structure provides a principled motivation for restart/windowing/forgetting rules that truncate stale history and thereby reduce both drift accumulation and variance inflation after regime changes.
@@
    Minimax lower bounds and the “inertia window” as an information-theoretic barrier for momentum. We establish minimax dynamic-regret lower bounds over a standard gradient-variation budget. The lower bound contains an explicit momentum-dependent term and formalizes an unavoidable inertia window: with large momentum, the method is provably “behind” a drifting optimizer for a nontrivial period, regardless of tuning. In the uniformly spread drift regime, our tuned upper bound matches the minimax rate and exhibits the sharp momentum penalty.
@@
    Nonstationary numerical experiments that corroborates the predicted regime split. We evaluate SGD, Heavy-Ball, and Nesterov on drifting quadratics, drifting linear/logistic regression, and a drifting teacher–student MLP. Across tasks, the results match the theory: momentum can help in near stationary/noise-dominated settings, while under genuine drift—especially with large $\beta$ or ill-conditioning—momentum becomes brittle and SGD is uniformly more robust.
@@ 1.1 Related Work
1.1.1 Momentum and Acceleration in Stochastic Optimization
Developing methods to accelerate the convergence of SGD has been studied extensively. Momentum methods, such as Polyak’s heavy-ball (HB) [POL63] and Nesterov acceleration (NAG) [NES83], are ubiquitous in deep learning and are widely believed to reduce gradient noise and accelerate training [SMD+13]. Adaptive variants such as Adagrad and Adam further modify the effective step-size online [DHS11, KB15, KB17, LH19, ZEI12, LK24], and the general research landscape of second-order and filtering based methods has been studied greatly [YAN23, YGS+21, OLL19, VUC18, DSC+22, ZRS+18, ZDT+21, ZJD+20, LLZ+25, GLP25, ZZL+19, IKS22, LJH+20a, JKR25, ZWU21, LXL19, GKS18, VMZ+25, YYC+25, YYC+25, DOZ16], with methods being developed specifically for optimizing the training of large neural networks as of late [LSY+25, HZJ+25, BWA+18]. It is widely known for a variety of strongly convex and nonconvex settings that SGD with momentum provides faster convergence and better generalization compared to SGD without momentum [RAC+24, SI24, LGY20b, CM20]. While it is relatively well-known that a fixed momentum rate is not as effective as adaptive momentum [SMD+13], addressing how to tune momentum optimally is an active area of research. In practice, it is common to fix the momentum parameter to a standard value (e.g., $\beta=0.9$) [KB17, CPM+17], or use a simple scheduler such as exponential momentum decay, cosine momentum, etc.
1.1.2 Limitations and Instabilities of Momentum
While popular, momentum does not always improve training [WMW+24, FWZ+23, DSL+24]. For online and adaptive learning, prior work suggests that the deterministic acceleration intuition does not directly transfer: in the slow-adaptation regime, momentum can be fundamentally equivalent to standard SGD with a re-scaled step-size, and can even degrade steady-state performance unless parameters are chosen carefully [YYS16]. A key challenge in nonstationary time series is that past gradients become stale: gradients computed under $\Pi_{t-k}$ can systematically point in the wrong direction for the current objective $G_{t}$. This motivates discounting or forgetting historical gradient information [ZWL20]. Furthermore, empirical evidence in online forecasting has pointed to SGD with momentum not being as stable as SGD without momentum, especially as learning rates grow [AZF19]. In [JJS25] they show that in high dimensional online settings, momentum can amplify noise and degrade the performance of SGD when step size is kept constant. In [DSL+24], they show that momentum degrades the stability of SGD when $\beta$ approaches $1$. Methods have been developed to tune momentum throughout the training process in a more stable manner, such as using local quadratic approximations [ZM19], combining different loss planes [TC25], momentum decay [CWL+22], or passive damping [LSZ+19]. Although research has been done on understanding the effects of momentum theoretically [JL22, LK24, GLZ+19, RKK18, QIA99, KS21], the noisy, time-dependent, non-stationary environment is understudied.
1.1.3 Optimization under Time-Dependent and Non-Stationary Environments
In [YYS16] they establish that in online stochastic optimization with constant step-sizes, momentum methods are essentially equivalent to standard SGD with a re-scaled learning rate at the steady-state. Their results indicate that momentum does not inherently lower the mean-square error (MSE) floor under persistent noise when the target is fixed. However, their analysis is primarily situated in the stationary regime; it does not account for the tracking lag and stability issues that arise when the underlying minimizer $\bm{\theta}_{t}^{\star}$ evolves as a non-stationary process. This is also supported in [KS21]. In [YPT17], they present a novel online gradient learning method to deal with time dependent data containing outliers by modifying the Adam Optimizer to suggest which data points are outliers. In [GTG+24], they do a soft parameter reset for neural network parameters in the non-stationary regime modeled as an Ornstein-Uhlenbeck process with an adaptive drift parameter that pulls weights toward a prior distribution. In [CDH23] they give non-asymptotic efficiency estimates for the tracking error of the proximal stochastic gradient method under time drift. In [SZZ26] they establish optimal estimation and regret bounds for SGD under temporally dependent data. [VID25] studies the stability and convergence of momentum-based methods where the objective function is allowed to vary with time. They do not address non-stationarity explicitly nor do they prove any bounds on the tracking error. In [CZP19], they study online stochastic optimization under time-varying data distributions and derive dynamic regret bounds for SGD. In [ZZZ+24] they study online convex optimization in non-stationary environments and develop problem-dependent dynamic regret bounds that adapt to the variation of the optimal solution and the loss sequence over time.
@@ 2 Preliminaries
2.1 Notations
We will denote scalars and vectors by lowercase and bold lowercase letters, respectively, and denote matrices by uppercase boldface letters. We use calligraphic letters to denote sets, operators, or $\sigma$-algebras. For a vector $\mathbb{x}=(x_{1},\dots,x_{d})\in\mathbb{R}^{d}$, we denote the $\ell_{2}$ norm by $\|\mathbb{x}\|$. We also denote the inner product between two vectors $\mathbb{x},\mathbb{x}^{\prime}\in\mathbb{R}^{d}$ as $\langle\mathbb{x},\mathbb{x}^{\prime}\rangle$. The transpose of a matrix $\mathbb{A}$ is indicated by $\mathbb{A}^{\top}$.
Given two sequences $\left\{a_{m}\right\}$ and $\left\{b_{m}\right\}$, we write $a_{m}=\mathcal{O}(b_{m})$ if there exists a positive constant $0<C<+\infty$ such that $a_{m}\leq Cb_{m}$, $a_{m}=\Theta(b_{m})$ if there exists two positive constants $C_{1},C_{2}>0$ such that $a_{m}\leq C_{1}b_{m}$ and $b_{m}\leq C_{2}a_{m}$, and $a_{m}=\Omega(b_{m})$ if there exists a positive constant $0<C<+\infty$ such that $a_{m}\geq Cb_{m}$. We write $a_{m}\lesssim b_{m}$ if there exists a constant $C>0$ that may depend only on fixed absolute constants (e.g., universal numerical constants) but not on $m$ or the varying problem parameters such that $a_{m}\leq C\,b_{m}$. We define $a_{m}\gtrsim b_{m}$ and $a_{m}\asymp b_{m}$ analogously.
For a function $f(\mathbb{x}):\mathbb{R}^{d}\rightarrow\mathbb{R}$, we denote its gradient as $\nabla f(\mathbb{x})$. The minimum value of $f(\mathbb{x})$ is denoted as $\min f$ and the minimizer of $f(\mathbb{x})$ is denoted as $\mathbb{x}^{*}$. We use $\mathbb{E}[\cdot]$ to denote the expectation with respect to the underlying probability measure. We also denote $\mathcal{F}_{t}=\sigma(X_{0},X_{1},\dots,X_{t})$ to denote the natural filtration i.e. the $\sigma$-algebra generated by random variables $X_{0},\cdots,X_{t}$.
2.2 Orlicz norm and conditional Orlicz norms
Let $X$ be a real-valued random variable and let $\alpha\geq 1$. The $\Psi_{\alpha}$–Orlicz norm is
$\|X\|_{\Psi_{\alpha}}\;:=\;\inf\Big\{u>0:\ \mathbb{E}\exp\!\big(|X/u|^{\alpha}\big)\leq 2\Big\}.$
Let $\mathcal{F}$ be a $\sigma$-field. A random variable $K_{\mathcal{F}}$ is said to be $\mathcal{F}$-measurable if $K_{\mathcal{F}}$ is measurable with respect to $\mathcal{F}$ (equivalently, $K_{\mathcal{F}}\in\mathcal{F}$).
Definition 2.1 (Conditional Orlicz norm).
Let $\mathcal{F}$ be a $\sigma$-field and let $K_{\mathcal{F}}>0$ be an $\mathcal{F}$-measurable random variable. We say that $X$ satisfies $\|X\mid\mathcal{F}\|_{\Psi_{\alpha}}\leq K_{\mathcal{F}}$ if and only if
$\mathbb{E}\!\left[\exp\!\Big(\big|X/K_{\mathcal{F}}\big|^{\alpha}\Big)\,\middle|\,\mathcal{F}\right]\leq 2,\qquad\text{a.s.}$ (2.1)
Note that the conditional expectation in (2.1) is itself an $\mathcal{F}$-measurable random variable. For a random vector $\mathbf{X}\in\mathbb{R}^{d}$, define
$\|\mathbf{X}\|_{\Psi_{\alpha}}\;:=\;\sup_{\mathbf{u}\in\mathbb{S}^{d-1}}\ \|\mathbf{u}^{\top}\mathbf{X}\|_{\Psi_{\alpha}},\qquad\mathbb{S}^{d-1}:=\{\mathbf{u}\in\mathbb{R}^{d}:\|\mathbf{u}\|_{2}=1\}.$
Definition 2.2 (Conditional Orlicz norm (vector)).
Let $\mathcal{F}$ be a $\sigma$-field and let $K_{\mathcal{F}}>0$ be $\mathcal{F}$-measurable. We say that $\mathbf{X}$ satisfies $\|\mathbf{X}\mid\mathcal{F}\|_{\Psi_{\alpha}}\leq K_{\mathcal{F}}$ if and only if
$\sup_{\mathbf{u}\in\mathbb{S}^{d-1},\mathbf{u}\in\mathcal{F}}\mathbb{E}\!\left[\exp\!\Big(\big|\mathbf{u}^{\top}\mathbf{X}/K_{\mathcal{F}}\big|^{\alpha}\Big)\,\middle|\,\mathcal{F}\right]\leq 2,\qquad\text{a.s.}$ (2.2)
For a random matrix $\mathbf{X}\in\mathbb{R}^{d_{1}\times d_{2}}$, define
$\|\mathbf{X}\|_{\Psi_{\alpha}}\;:=\;\sup_{\mathbf{u}\in\mathbb{S}^{d_{1}-1}}\ \sup_{\mathbf{v}\in\mathbb{S}^{d_{2}-1}}\ \|\mathbf{u}^{\top}\mathbf{X}\mathbf{v}\|_{\Psi_{\alpha}},$
with conditional versions defined analogously (conditioning on $\mathcal{F}$ and allowing $\mathcal{F}$-measurable unit vectors in the suprema). These norms are well-defined and satisfy the usual norm properties (see Section˜D.1 for more details).

### v3

@@ Abstract
In this paper, we provide a comprehensive theoretical analysis of Stochastic Gradient Descent (SGD) and its momentum variants (Polyak Heavy-Ball and Nesterov) for tracking time-varying optima under strong convexity and smoothness. Our finite-time bounds reveal a sharp decomposition of tracking error into transient, noise-induced, and drift-induced components. This decomposition exposes a fundamental trade-off: while momentum is often used as a gradient-smoothing heuristic, under distribution shift it incurs an explicit drift-amplification penalty that diverges as the momentum parameter $\beta$ approaches 1, yielding systematic tracking lag. We complement these upper bounds with minimax lower bounds under gradient-variation constraints, proving this momentum-induced tracking penalty is not an analytical artifact but an information-theoretic barrier: in drift-dominated regimes, momentum is unavoidably worse because stale-gradient averaging forces systematic lag. Our results provide theoretical grounding for the empirical instability of momentum in dynamic settings and precisely delineate regime boundaries where vanilla SGD provably outperforms its accelerated counterparts.
Nonstationary stochastic optimization,Momentum methods,Tracking error bounds,Dynamic regret under gradient-variation budgets,Minimax lower bounds / information-theoretic limits
@@ 1 Introduction
Consider the optimization problem posed by a strongly convex objective function $G_{t}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ defined as follows:
$\bm{\theta}_{t}^{\star}\in\operatorname*{arg\,min}_{\bm{\theta}\in\mathbb{R}^{d}}G_{t}(\bm{\theta}),\;\;G_{t}(\bm{\theta})=\mathbb{E}_{X_{t}\sim\Pi_{t}}[g(\bm{\theta},X_{t})]$ (Opt)
Here $g(\bm{\theta},X_{t})$ is a noise perturbed measurement of $G_{t}(\bm{\theta})$ and $X_{t}$ is a random variable sampled from a time-varying distribution $\Pi_{t}$. These settings naturally connect to classical stochastic tracking in signal processing (Kushner and Yin, 1997; Sayed, 2003) and concept drift in online learning (Hazan, 2016; Cutler et al., 2021), where the goal is not to converge to a single static solution, but to track an unknown, time-varying sequence of minimizers as closely as possible.
The primary algorithmic tool for this tracking task, provided that $g$ is differentiable, is Stochastic Gradient Descent (SGD) (Robbins and Monro, 1951). Due to its simplicity, scalability to big data, and robustness to noise and uncertainty, stochastic gradient methods have become popular in large-scale optimization, machine learning, and data mining applications (Zhang, 2004; Cevher et al., 2014; Szegedy et al., 2015).
Developing methods to accelerate the convergence of SGD has been studied extensively. Momentum methods, such as Polyak’s Heavy-Ball (HB) (Polyak, 1963) and Nesterov acceleration (NAG) (Nesterov, 1983), are ubiquitous in deep learning and are widely believed to reduce gradient noise and accelerate training (Sutskever et al., 2013). For a variety of strongly convex and nonconvex settings, it is widely known SGD with momentum (SGDM) provides faster convergence and better generalization compared to SGD without momentum (Cutkosky and Mehta, 2020; Liu et al., 2020b; Sato and Iiduka, 2024; Ramezani-Kebrya et al., 2024). However, this deterministic intuition becomes fragile in nonstationary regimes: when the data distribution $\Pi_{t}$ drifts, past gradients may become systematically stale, so the same temporal averaging that suppresses noise can induce tracking lag and even instability, as observed empirically in online forecasting and suggested by prior steady-state analyses in the slow-adaptation regime.
@@
When does momentum help under distribution shift and when does it provably hurt for strongly convex, smooth functions $G_{t}(\bm{\theta})$?
@@
    Finite-time tracking with an explicit momentum penalty. We prove finite-time tracking bounds that separate three terms: initialization decay, a noise floor, and drift-induced tracking lag. Extending to Heavy-Ball and Nesterov yields the same decomposition, but with explicit momentum penalties—large $\beta$ amplifies initialization effects, inflates the noise floor, and imposes stricter stability constraints, particularly in ill-conditioned settings.
@@
    Time-resolved high-probability bounds and a drift–noise coupling. Unlike prior results based on a single worst-case drift parameter, our guarantees are time-resolved, with exponential forgetting ensuring only recent drift matters. Our analysis also reveals a drift–noise interaction whereby tracking mismatch amplifies stochastic fluctuations, motivating restart and windowing schemes that discard stale history.
@@
    Minimax lower bounds and an information-theoretic inertia window. We establish minimax dynamic-regret lower bounds under a gradient-variation budget with an explicit $\beta$-dependent term, yielding an unavoidable inertia window: for large momentum, SGDM must lag a drifting optimizer for a nontrivial period regardless of tuning. In the uniformly spread drift regime, our tuned upper bound matches the minimax rate and exhibits the sharp momentum penalty.
@@
    Experiments corroborating the regime split. On drifting quadratics, drifting linear/logistic regression, and a drifting teacher–student MLP, our results match the theory: increasing nonstationarity, momentum, or ill-conditioning reliably worsens HB/NAG tracking, consistent with an inertia-limited lag under drift, while SGD remains comparatively robust across noise levels.
@@ 1.1 Related Work
Momentum methods—including Polyak’s Heavy-Ball (Polyak, 1963) and Nesterov acceleration (Nesterov, 1983)—are ubiquitous in deep learning, where they are widely believed to reduce gradient noise and accelerate training (Sutskever et al., 2013). Indeed, for a variety of strongly convex and nonconvex settings, SGD with momentum provides faster convergence and better generalization than vanilla SGD (Cutkosky and Mehta, 2020; Liu et al., 2020b; Sato and Iiduka, 2024; Ramezani-Kebrya et al., 2024). However, momentum does not universally improve training and can degrade performance in online and adaptive settings (Fu et al., 2023; Wang et al., 2024; Deng et al., 2024). In nonstationary environments, past gradients become stale and may point in the wrong direction, motivating forgetting or discounting of historical information (Zhang et al., 2020). Empirically, momentum can also reduce stability at large learning rates (Aydore et al., 2019). Recent theoretical work reinforces these concerns: in high-dimensional or slow-adaptation regimes, momentum can amplify noise, undermine stability, or collapse to SGD with a rescaled step size (Yuan et al., 2016; Deng et al., 2024; Jagannath et al., 2025).
These findings connect to a broader literature on optimization in time-dependent and nonstationary environments. In stationary settings with constant stepsizes, momentum is essentially equivalent at steady state to SGD with a rescaled learning rate and does not reduce the MSE floor (Yuan et al., 2016; Kovachki and Stuart, 2021). Other work addresses nonstationarity from complementary perspectives: robust and adaptive methods for corrupted or drifting data (Yang et al., 2017; Galashov et al., 2024); non-asymptotic tracking and regret guarantees under time drift or dependent data (Cao et al., 2019; Cutler et al., 2023; Zhao et al., 2024; Shen et al., 2026); and stability analyses of momentum for time-varying objectives (Vidyasagar, 2025). Despite this progress, the interaction between momentum and nonstationarity remains significantly understudied—a gap this paper aims to address. To provide the necessary context, we survey the relevant literature on momentum and acceleration in nonstationary stochastic optimization more extensively in Appendix A.
@@ 2 Preliminaries
We first summarize the notation used throughout the paper. Scalars, vectors, and matrices are denoted by lowercase, bold lowercase, and bold uppercase letters, respectively; calligraphic letters denote sets, operators, or $\sigma$-algebras. For $\mathbb{x}\in\mathbb{R}^{d}$, $\|\mathbb{x}\|$ is the $\ell_{2}$ norm and $\langle\mathbb{x},\mathbb{x}^{\prime}\rangle$ the inner product. We write $a_{m}=\mathcal{O}(b_{m})$ if $a_{m}\leq Cb_{m}$, $a_{m}=\Omega(b_{m})$ if $a_{m}\geq Cb_{m}$, and $a_{m}=\Theta(b_{m})$ if both hold, for some $C>0$. The notation $a_{m}\lesssim b_{m}$ (resp. $\gtrsim$, $\asymp$) indicates inequality up to constants independent of $m$ and problem parameters. For $f:\mathbb{R}^{d}\to\mathbb{R}$, we write $\nabla f$ for the gradient, $\min f$ for the minimum value, and $\mathbb{x}^{*}$ for the minimizer. We use $\mathbb{E}[\cdot]$ for expectation and $\mathcal{F}_{t}=\sigma(X_{0},\ldots,X_{t})$ for the natural filtration.
We next introduce conditional $\Psi_{\alpha}$–Orlicz norms, which we will use throughout our analysis to control random quantities in a filtration-adapted (i.e., history-dependent) manner. Fix $\alpha\geq 1$. For a real-valued random variable $X$, recall the (unconditional) $\Psi_{\alpha}$–Orlicz norm
$\|X\|_{\Psi_{\alpha}}:=\inf\Big\{u>0:\ \mathbb{E}\exp\big((|X|/u)^{\alpha}\big)\leq 2\Big\}.$
Given a $\sigma$-algebra $\mathcal{F}$, we write $\|X\|_{\Psi_{\alpha}\mid\mathcal{F}}\leq K_{\mathcal{F}}$ for some $\mathcal{F}$-measurable $K_{\mathcal{F}}>0$ if $\mathbb{E}\!\left[\exp\!\left(\big(|X|/K_{\mathcal{F}}\big)^{\alpha}\right)\middle|\mathcal{F}\right]\leq 2$. Equivalently, one may define for $u$
$\|X\|_{\Psi_{\alpha}\mid\mathcal{F}}:=\inf\Big\{u>0:\ \mathbb{E}\!\left[\exp\big((|X|/u)^{\alpha}\big)\mid\mathcal{F}\right]\leq 2\ \Big\}$
provided $u$ is $\mathcal{F}$ measurable. Vector and matrix conditional Orlicz norms are defined analogously to their unconditional counterparts by taking suprema over one-dimensional projections (see Appendix F.1 for details).